OpenStudy (anonymous):

sum_(k=1)^6-3(2)^k-1

OpenStudy (anonymous):

\[\sum_{k=1}^6{3(-2)^k-1}\] The -1 is up with the k. Find the indicated sum of the geometric series.

OpenStudy (anonymous):

\[\sum_{k=1}^6{3(-2)^{k-1}}\]

OpenStudy (anonymous):

Find the indicated sum of the geometric series.

OpenStudy (anonymous):

with so few terms i would just add them

OpenStudy (anonymous):

How so? And I messed up. It is -3(2). Sorry

OpenStudy (anonymous):

\[\sum_{k=1}^6{3(-2)^{k-1}}=3\sum_{k=1}^6{(-2)^{k-1}}\] \[=3(1-2+4-8+16-32)\]

OpenStudy (anonymous):

\[\sum_{k=1}^6-3(2)^{k-1}\]

OpenStudy (anonymous):

like that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\sum_{k=1}^6-3(2)^{k-1}=-3\sum_{k=1}^6(2)^{k-1}\]

OpenStudy (anonymous):

\[-3\sum_{k=1}^6(2)^{k-1}=-3(1+2+4+8+16+32)=-3(63)\] if my arithmetic is correct

OpenStudy (anonymous):

Yes! Thank you so much satellite73!!!! :)

OpenStudy (anonymous):

yw